Fermi National Laboratory


Accelerator Update

Wednesday February 18
The day shift began with Operations monitoring store 3240, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8 AM, Tevatron (TeV) experts began an end-of-store study.

Switchyard experts began a study period at 8:47 AM.

Operations terminated store 3240 at 11:10 AM.

At 11:26 AM, Operations made the Collider Detector at Fermilab (CDF) and D-zero (D0) ready for access.

At 12:02 PM, Operations had the TeV ready for a two-hour access to conduct vacuum work.

CDF came out of access at 12:46 PM and D0 came out at 1:45 PM.

The Main Injector (MI), Recycler, and MiniBooNE had all started a study period by 4 PM.

Switchyard studies ended at 6:51 PM.

Operations put the TeV into shot setup and at 10:07 PM established store 3245. Its initial luminosity was 63.25E30. That is a new luminosity record! Pbar resumed stacking.

Thursday February 19
The midnight shift began with Operations monitoring store 3245, with Pbar Stacking, and with MiniBooNE taking beam.

At 2 AM, Operations began stashing antiprotons to the Recycler.

Stashing ended and Pbar resumed stacking at 3:13 AM. But the stack rate was poor. Operators replaced a power supply (D:V401) at 4:22 AM, which fixed the problem.

At 5 AM, the Time Line Generator (TLG) failed. Operations rebooted the system and re-established beam at 5:20 AM.

Switchyard studies began at 8:29 AM.

Operations began stashing antiprotons to the Recycler at 11:11 AM. Stacking resumed at 11:40 AM.

At 1:28 PM, a Linac expert conducted a thirty-minute study.

At 1:36 PM, a Booster expert conducted a two-hour collimator study.

Friday February 20
The midnight shift began with Operations monitoring store 3245 and the Recycler stash, with Pbar Stacking, and with MiniBooNE taking beam.

At 2:22 AM, Operations received an alarm for a LCW leak in the Booster tunnel. Operators turned off beam and prepared Booster for an access. At 3:45 AM, they reported closing valves around a bad hose. Operations called in water technicians to make repairs and RF techs to change the power amplifier for BRF2.

Water techs finished their Booster hose repair at 6 AM.

RF techs completed their work in the tunnel on BRF2 at 7 AM. They then began work on equipment outside of the tunnel.

Operations had the Booster ready to take beam by 7:42 AM.

By 7:57 AM, Operations had established beam to MiniBooNE and Pbar had resumed stacking.

Today's Plans
The plans for today include letting Booster experts tune their beam and letting TeV experts conduct a synclight (beam position) study. After that, Operations will terminate the store and then put the TeV into shot setup later this morning.

Operations will use the Recycler stash to establish the next TeV store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/20/2004   email Fermilab